OpenStudy (anonymous):

Can you help me solve this integral and explain how it's done ??? 8x/(-4x^4+12x^2+27)^1/2 dx

OpenStudy (anonymous):

\[\int\frac{8x}{\sqrt{-4x^4+12x^2+27}}~dx~~?\] Complete the square in the denominator. \[\begin{align*} -4x^4+12x^2+27&=-4\left(x^4-3x^2-\frac{27}{4}\right)\\\\ &=-4\left(x^4-3x^2+\frac{9}{4}-\frac{36}{4}\right)\\\\ &=-4\left(\left(x^2-\frac{3}{2}\right)^2-\frac{36}{4}\right)\\\\ &=36-4\left(x^2-\frac{3}{2}\right)^2 \end{align*}\] Use a substitution, \(u=x^2-\dfrac{3}{2}\), then \(\dfrac{1}{2}~du=x~dx\). \[4\int\frac{du}{\sqrt{36-u^2}}\] Make another substitution, \(u=6\sin t\), so that \(du=6\cos t~dt\). \[24\int\frac{\cos t}{\sqrt{36-36\sin^2t}}~dt\\ 4\int dt\]

OpenStudy (anonymous):

sorry, ammmm why in the second substitution u= 6sinx & du = 6cosx dx i can't understand xD

OpenStudy (anonymous):

For the \(u\)-integral \[4\int\frac{du}{\sqrt{36-u^2}}\] you need a substitution that can work with the otherwise difficult-to-integrate expression in the denominator. For expressions of this form, we use some useful trig identities. Recall the Pythagorean identity:\[\sin^2x+\cos^2x=1\] which gives \(1-\sin^2x=\cos^2x\). Note the similarity between the left side of this identity and the denominator - we have a difference of (not necessarily perfect) squares. For this reason, we'll let \(u=6\sin t\), so that \(du=6\cos t~dt\). (Nothing fancy here, just taking thing the differentials.) Then you substitute as you normally would: \[\begin{align*} 4\int\frac{\color{blue}{du}}{\sqrt{36-\color{red}u^2}}&=4\int\frac{\color{blue}{6\cos t}}{\sqrt{36-\color{red}{(6\sin t^2)}}}~\color{blue}{dt}\\\\ &=24\int\frac{\cos t}{\sqrt{36-36\sin^2t}}~dt\\\\ &=\frac{24}{\sqrt{36}}\int\frac{\cos t}{\sqrt{1-\sin^2t}}~dt\\\\ &=\frac{24}{6}\int\frac{\cos t}{\sqrt{\cos^2t}}~dt\\\\ &=4\int\frac{\cos t}{\cos t}~dt\\\\ &=4\int dt \end{align*}\]